# Excitation Spectrum and Spin Hamiltonian of the Frustrated Quantum Ising Magnet PrBWO

### v1

@@ III.4 Model Hamiltonian
Due to the large separation of energy scales between crystal field and exchange interactions, we can construct a low-energy effective Hamiltonian retaining only the quasi-doublet states. The only relevant component from the CEF Hamiltonian will be the splitting between them. It is well established [wangCollectiveExcitationsMagnetic1968, savaryDisorderInducedQuantumSpin2017, chenIntrinsicTransverseField2019] that an energy splitting between two CEF singlets can be mapped exactly into a transverse magnetic field acting on a corresponding doublet. The resulting single-ion Hamiltonian can be written in terms of local pseudospins as
where is the projector onto the local Ising axis. Throughout this paper, the -notation always refers to the local basis of the corresponding magnetic ion.
Moreover, the projected two-ion interactions are also severely restricted by the nature of the single-ion ground state. Due to the strict Ising anisotropy, the transverse angular momentum components vanish between the quasi-doublet states. This immediately results in an Ising form in the local basis for bilinear exchange coupling , be it from superexchange or dipolar coupling. Even the higher multipolar interactions typical in rare earths can be ignored here, as they are restricted to rank by the maximal angular momentum transferred in each step of the superexchange process and as such cannot connect the leading ground state components [rauMagnitudeQuantumEffects2015, iwaharaExchangeInteractionMultiplets2015].
@@
where is the longitudinal -factor and the weak dipolar interactions of order K are simply absorbed into the effective Ising exchange parameters . Potential non-Ising like deviations from this Hamiltonian are constrained by the small mixing of lower angular momentum states into our refined ground state wavefunctions. As such, PrBWO turns out to be an excellent realization of the transverse field Ising model (TFIM) Hamiltonian.
@@ III.5 Heat Capacity
The zero field heat capacity of PrBWO is shown in Figure 5, where the magnetic contribution is extracted by subtracting the nonmagnetic LaBWO reference data. We observe a broad hump at high temperatures associated with a depopulation of excited crystal field states, as well as another Schottky-like peak centered around 10 K, which matches the energy scale of the meV ( K) level. Comparing these data to the single-ion contribution expected from our CEF model, is nicely reproduced down to K, with clear deviations below especially visible on a log scale in Figure 5(b). A small part of the expected 1.7 meV Schottky feature seems to be shifted towards lower temperatures, where we observe a power-law-like decay with instead of the activated behavior of a lone Pr ion, with perhaps a tiny upturn below 0.2 K. Importantly there is no sign of long-range order down to K, consistent with previous reports [zengLocalEvidenceCollective2021]. This redistribution of spectral weight could be associated with magnetic correlations, the onset roughly matching the mean field exchange energy meV extracted above. Still, the close match to single-ion calculations down to might suggest interactions are significantly weaker than the crystal field splitting, making a strongly correlated spin-liquid like state rather unlikely. Although the gapless behavior at low temperatures is seemingly inconsistent with a singlet ground state system, it can arise naturally in the presence of structural disorder as we will show in Sec. III.8, broadening the mode at into a distribution of crystal field gaps. The change in magnetic entropy seen in Figure 5(d) is extracted through numerical integration of the heat capacity data. As expected for a well isolated quasi-doublet system, it closely approaches for temperatures , justifying the simplified model Hamiltonian derived in Sec. III.4.
@@
Towards lower temperatures an additional broad peak-like feature appears at the scale of K, showing a linear field dependence (see Figure 6(c)). However, the associated -factor is much too small to be of electronic origin, instead matching calculated for hyperfine coupled nuclear moments [bleaneyEnhancedNuclearMagnetism1973]. In fact, the onset temperature of this feature agrees nicely with that of the calculated nuclear Schottky contribution, as shown exemplarily for the 14 T data in Figure 6(a). Nevertheless, the magnetic entropy released corresponds only to of and the feature becomes suppressed slightly below the onset temperature of .
Although a drastically reduced nuclear heat capacity might seem puzzling, there are several potential mechanisms for bringing this about. For example, it is known that electronic spin fluctuations can lead to a reduction in the nuclear Schottky contribution [bertinEffectiveHyperfineTemperature2002]. The corresponding fluctuations in the hyperfine field can diminish the population differences between the Pr Zeeman levels, resulting in a suppressed , seen e.g. in Pr-based pyrochlores [maclaughlinUnstableSpiniceOrder2015, martinDisorderQuantumSpin2017a]. Another possibility would be a partial decoupling of the nuclear degrees of freedom due to a lack of phonons (or magnons) at these temperatures [abragamPrinciplesNuclearMagnetism, isonoSpinlatticeDecouplingTriangularlattice2018a]. The nuclear spin system only couples indirectly to the external heat bath through excitations of the lattice/electronic spins, so without an abundance of such low energy modes we cannot probe efficiently the nuclear heat capacity. A similar decoupling phenomenon was observed in the sister material NdBWO above the pseudospin saturation [flavianMagneticPhaseDiagram2023a].
In PrBWO, the purported nuclear contribution is suppressed immediately below the onset temperature of the calculated , which heavily favours the latter scenario. For fast fluctuating electronic spins, the peak position should be preserved, while a decoupling of nuclear spins would become effective for , meaning that only the high temperature tail of the nuclear heat capacity can be observed. Towards lower fields, the hyperfine splitting decreases significantly, further suppressing until it is no longer visible below T.
@@ III.6 Ultrasound
Figure 7 displays the relative change in the sound velocity of PrBWO as a function of magnetic field. We probe the longitudinal mode in this experiment, propagating with km/s. In agreement with the heat capacity data, we do not detect any signs of phase transitions down to mK even for fields exceeding 40 T, suggesting that PrBWO retains a quantum disordered ground state in the entire phase space. Instead, we observe a small softening at low fields followed by a strong quasi-linear increase in sound velocity. Both the earlier crossover between these two regimes and the overall stronger magnetoelastic response for fields along are linked to the easy-axis nature of PrBWO, explained mostly by a difference in -factors. Furthermore, qualitatively follows the single magnon gap (see Figure 6(b)), where the crossover appears when the crystal field splitting acts on the same scale as the Zeeman energy . In fact, these observations are not surprising, as the dominant contribution to the magnetoelastic energy density in rare earths typically originates from a linear single-ion strain interaction [luthiPhysicalAcousticsSolid1967]. Lattice vibrations modulate the crystal electric field, which directly determines the orbital state of each ion, resulting in a magnetoelastic response dictated by single-ion properties. Though the overall scale of of course depends on the field orientation, the sign of the effect does not. Because there is no transverse -factor, fields can only couple to the local spin component, resulting in an analogous deformation to the charge density independent of field direction. We note that in principle, the phase sensitive detection method utilized here responds also to magnetostrictive changes in the sample length along the propagation direction, i.e. we measure the phase shift as denoted in Figure 7. However, barring any (magneto-)structural phase transitions or incipient soft modes thereof, the change in lattice spacing should be orders of magnitude weaker than the present effects and will therefore be ignored here.
@@
Perhaps more interesting is the significant temperature dependence at low fields shown in the inset of Figure 7. There is a rapid hardening with increasing temperature at zero field, which becomes heavily suppressed at fields as low as T, gradually giving way to -independent behavior. All data are taken below , so effects of CEF population should be negligible and these features more likely result from a change in the spin correlations. But even though the lattice hardening at zero field can be easily rationalized as a weakening of magnetic correlations, it is unclear how its sudden suppression at extremely low fields could be explained in this scenario. Indeed, there is no obvious energy scale associated with this suppression - both Weiss temperature and crystal field gap are significantly larger - and neither susceptibility nor magnetization exhibit significant features at such low fields, making this behavior somewhat anomalous.
@@ III.7 Muon Spin Relaxation
As a complement to thermodynamic probes, the ground state magnetism in PrBWO is characterized by means of muon spin relaxation. The asymmetry spectra of the zero-field SR data at selected temperatures are displayed in Figure 8(a). Contrary to expectations, the low temperature data show a clear dip at s followed by the partial recovery of a weakly decaying tail, amounting to of the initial polarization. The signal remains temperature independent below K, whereas at higher temperatures the dip weakens and shifts to later times, gradually giving way to paramagnetic relaxation above K. A Kubo-Toyabe-like (KT) -tail provides strong evidence for quasi-static magnetism at the SR timescale [leeMuonScienceMuons2017], whereas the lack of further oscillations implies a frozen disordered state instead of collective long-range order.
@@
With this model, we can successfully account for the zero field relaxation in the entire temperature range 30 mK 200 K - the resulting fit parameters are provided in Figure 9. The width of the internal field distribution remains constant below 5 K before smoothly starting to decrease, only tending towards zero at the highest measured temperatures. Both the relaxation rate and the KT-exponent show a broad hump around 50 K, consistent with a slow crossover towards paramagnetic relaxation at high temperatures. Apart of that, the only prominent feature in the data is a step-like change seen around K, present in all fit parameters except . Especially the step in is directly visible in Figure 8(a) as a shift of the late-time asymmetry tail between the 1 K and 2.5 K curves. In fact, the same feature remains present in a T longitudinal field (see Figure 8(b)), even though the weak internal fields must be fully decoupled in this regime, suggesting a dynamical origin for the anomaly. We can explain this behavior by invoking the existence of several stopping sites: The residual dynamics in some fraction of the sample slows down at low temperatures until the tail relaxation of that component approaches zero within our time window. This brings about an increase in the “background” fraction , as well as a change in the effective relaxation rate . It is also consistent with the appearance of a stretching exponent for the dynamical relaxation, which is expected in the presence of several different relaxation times. The only magnetic energy scale roughly matching this K feature would be the that of the two-ion interactions meV, indicating that it could be related to the onset of some residual AFM correlations, which quench/slow down the dynamics in a fraction of the sample.
@@
In order to pinpoint the nature of the weak quasi-static magnetism, we compare the mT observed at low temperatures to typical internal field values expected in different scenarios. For instance, similar Pr-based materials that order magnetically typically exhibit a strong T dipolar field at the muon site [AnandPRB1072023], roughly two orders of magnitude larger than seen in experiment. This clearly contradicts the possibility of a glassy ground state, where the SR relaxation would be dominated by fully frozen electronic moments . Moreover, the internal field distribution is also atypical of nuclear magnetism (the Pr nuclei carry a spin ), which usually leads to a temperature-independent width in the range of mT - seemingly ruling out nuclear moments as a source of the KT-like behavior. But as we are dealing with a non-Kramers system endowed with strong hyperfine coupling and a low-lying CEF level, there is a well known mechanism for amplifying the effective nuclear moments, aptly named enhanced nuclear magnetism [bleaneyEnhancedNuclearMagnetism1973, blundellQuantumMuon2023]. Due to the large Van-Vleck susceptibility, the nuclear spins can partially polarize the electronic shell, giving rise to “dressed” nuclear moments
@@ III.8 Inelastic Neutron Scattering
Distance (Å) |
Value (meV) |
Value (meV) |
1.62(1) |
1.62(1) |
3.96 |
0.30(1) |
0.0068(2) |
3.96 |
-0.11(1) |
-0.0026(2) |
4.26 |
0.06(1) |
0.0053(8) |
The inelastic neutron spectra collected at AMATERAS are summarized in Figure 10 and Figure 11. These are raw data without background subtraction, taken at mK. The spectrum in PrBWO seems to be composed of two kinds of excitations, featuring a coexistence of sharp spin-wave-like modes and broad continuum excitations.
@@
In order to analyze the observed spectrum, we fit a number of constant- cuts to a combination of one Gaussian-broadened (continuum) and several resolution-limited (spin waves) modes, extracting almost 600 dispersion points across the entire Brillouin zone. These data are analyzed by means of linear spin-wave theory (LSWT) calculations using the SUNNY software package [sunny_git]. We perform a global LSWT fit to the observed mode energies, based on the two-level Hamiltonian Eq. (5) derived above. In accordance with our thermodynamic data, the ground state is assumed to be quantum disordered in these calculations, corresponding to a fully polarized state in the pseudospin language of our transverse-field Ising model, where each spin points along its local -axis. Therefore, LSWT is expected to be an excellent approximation and should allow for an accurate estimation of the “true” (non-renormalized) exchange constants [coldeaDirectMeasurementSpin2002].
@@
The calculated intensities can be improved significantly by using the full Hamiltonian
i.e. working with coherent states [dahlbomRenormalizedClassicalTheory2023] instead of dipoles polarized along the transverse field direction. In this model, the crystal field parameters are rescaled globally by a factor from their values in Table 5, which allows us to fine tune the observed gap energy while maintaining the anisotropies refined in bulk. The exchanges also have to be rescaled as
@@
to account for the change in spin length, where is the inelastic matrix element connecting our quasi-doublet states. The geometrical factor reflects the fact that we are using Heisenberg couplings in the full model; projecting down to the low-energy degrees of freedom will anyways render void all exchange components not along the local Ising axes. The appropriately rescaled exchange parameters are also listed in Table 3. These are fully equivalent to the transverse-Ising exchanges, producing the exact same dispersion, but give a quantitative agreement on the relative intensity distribution. The relatively stronger simply reflects the difference in angle between the easy axes of first and second neighbors.
@@
Finally, we attempt to empirically model the broad continuum excitations. Since all spin-wave branches are accounted for across the Brillouin zone and the thermodynamic probes rule out the presence of further accidentally degenerate CEF modes, it is unclear how any scenario of fractionalized excitations could give rise to the continuum seen in PrBWO. Instead, the fact that the latter seems to mimic the spin-wave dispersion favors a scenario of structural disorder. Due to the lack of time reversal symmetry in Pr, the splitting between our quasi-doublet states is highly sensitive to changes in the local crystal structure [martinDisorderQuantumSpin2017a, liSpinDynamicsGriffiths2021]. As such, any local distortion due to e.g. strains, impurities or domain walls will slightly alter the CEF scheme, resulting in a distribution of gaps . As we observe both broad and sharp contributions in our spectra, we have to assume that some ions are sufficiently removed from such impurity sites to give rise to the resolution limited spin-waves, while the remaining ones possess a random distribution of gaps, accounting for the continuum. An investigation of the local strain fields in PrBWO goes beyond the scope of this work, though we find that a simple Gaussian distribution is sufficient for modeling the spectrum.
@@
Indeed, we can successfully reproduce our inelastic data in this disorder scenario by adding a Gaussian-broadened copy of the calculated spin-wave spectrum with the appropriate spectral weight. To that end, we introduce three new global parameters: The fraction of affected magnetic sites, the Gaussian width meV reflecting the degree of randomness and the mean meV of the distribution (to obtain complete agreement with experiment we choose ). Presumably, quenched disorder would also induce some variation in the exchanges, but here we push all randomness onto the Gaussian width , in order not to over-parameterize our model. The calculated spectra obtained with these empirically optimized parameters are displayed in Figures 10(k,l,m,n,o) and 11(e,f,g,h). Our model can semi-quantitatively account for all experimental features, including the coexistence of resolution-limited and broadened modes, as well as the intensity distribution across the entire Brillouin zone.
@@
As a cross check, we calculate the single-ion heat capacity in the proposed disorder scenario, shown in Figure 5 as composed of Schottky terms with the distribution of gaps extracted from the neutron data. Unlike the “clean” single-ion calculation, this model closely approximates the measured curve in the entire temperature range with no free parameters, including the power-law like behavior at low temperatures. As seen in Figure 6, we can even reproduce the results at finite fields, although the small hump seen at low temperatures remains absent, consistent with our explanation in terms of partially decoupled nuclear degrees of freedom. This striking parameter-free agreement, together with the excellent fit of our inelastic neutron data, strongly justify our model assumptions, confirming that the magnetism of PrBWO should be understood in terms of dispersive crystal field excitons, with many sites affected by some form of structural disorder that locally tunes the average CEF scheme.

### v2

@@ III.4 Model Hamiltonian
Due to the large separation of energy scales between crystal field and exchange interactions, only the two lowest levels defined in Eq. (III.3) are populated at low temperatures and we can construct a low-energy effective Hamiltonian retaining only these quasi-doublet states. The only relevant component from the CEF Hamiltonian will be the splitting between them. It is well established [wangCollectiveExcitationsMagnetic1968, savaryDisorderInducedQuantumSpin2017, chenIntrinsicTransverseField2019] that an energy splitting between two CEF singlets can be mapped exactly into a transverse magnetic field acting on a corresponding doublet. The resulting single-ion Hamiltonian can be written in terms of local pseudospins as
where is the projector onto the local Ising axis. Throughout this paper, the -notation always refers to the local basis of the corresponding magnetic ion. Further details about the pseudospin projection are provided in Appendix. B.
Moreover, the projected two-ion interactions are also severely restricted by the nature of the single-ion ground state. Due to the strict Ising-like anisotropy, the transverse angular momentum components vanish between the quasi-doublet states. This immediately results in an Ising form in the local basis for bilinear exchange coupling , be it from superexchange or dipolar coupling. Even the higher multipolar interactions typical in rare earths can be ignored here, as they are restricted to rank by the maximal angular momentum transferred in each step of the superexchange process and as such cannot connect the leading ground state components [rauMagnitudeQuantumEffects2015, iwaharaExchangeInteractionMultiplets2015].
@@
where is the longitudinal -factor and the weak dipolar interactions of order K are simply absorbed into the effective Ising exchange parameters . Potential deviations from this Hamiltonian are constrained by the small mixing of lower angular momentum states into our refined ground state wavefunctions. As such, we can establish a direct mapping between the low-energy physics of our Pr3+ magnetic moments subject to the crystal electric field and that of Ising-like pseudospin moments in a strong transverse magnetic field - Pr3BWO9 turns out to be an excellent realization of the TFIM Hamiltonian. Throughout this paper, we will continue to employ both the CEF model and the TFIM, pointing out the particular framework used as we go along.
@@ III.5 Heat Capacity
The zero field heat capacity of Pr3BWO9 is shown in Figure 5, where the magnetic contribution is extracted by subtracting the nonmagnetic La3BWO9 reference data. We observe a broad hump at high temperatures associated with a depopulation of excited crystal field states, as well as another Schottky-like peak centered around 10 K, which matches the energy scale of the meV ( K) level. Comparing these data to the single-ion contribution expected from our CEF model, is nicely reproduced down to K, with clear deviations below especially visible on a log scale in Figure 5(b). A small part of the expected 1.7 meV Schottky feature seems to be shifted towards lower temperatures, where we observe a power-law-like decay with instead of the activated behavior of a lone Pr3+ ion, with perhaps a tiny upturn below 0.2 K. Importantly there is no sign of long-range order down to K, consistent with previous reports [zengLocalEvidenceCollective2021]. This redistribution of spectral weight could be associated with magnetic correlations, the onset roughly matching the mean field exchange energy meV extracted above. Still, the close match to single-ion calculations down to might suggest interactions are significantly weaker than the crystal field splitting, making a strongly correlated spin-liquid like state rather unlikely. Although the gapless behavior at low temperatures is seemingly inconsistent with a singlet ground state system, it can arise naturally in the presence of structural disorder as we will show in Sec. III.8, broadening the mode at into a distribution of crystal field gaps. The change in magnetic entropy seen in Figure 5(d) is extracted through numerical integration of the heat capacity data. As expected for a well isolated quasi-doublet system, it closely approaches for temperatures , justifying the simplified model Hamiltonian derived in Sec. III.4.
@@
Towards lower temperatures an additional broad feature appears at the scale of K. We estimate its peak position by subtracting the apparent power-law behavior observed at intermediate temperatures, revealing a linear field dependence (see Figure 6(c)). However, the associated -factor is much too small to be of electronic origin, instead matching 111Given a strong hyperfine coupling, the nuclear spins may slightly polarize the electronic shell, effectively resulting in an enhanced nuclear moment - i.e. an increased nuclear -factor. The enhancement factor can be calculated explicitly, see Sec. III.7 for details. calculated for hyperfine coupled nuclear moments [bleaneyEnhancedNuclearMagnetism1973]. In fact, the onset temperature of this feature agrees nicely with that of the calculated nuclear Schottky contribution, as shown exemplarily for the 14 T data in Figure 6(a). Nevertheless, the magnetic entropy released corresponds only to of and the feature becomes suppressed slightly below the onset temperature of .
Although a drastically reduced nuclear heat capacity might seem puzzling, there are several potential mechanisms for bringing this about. For example, it is known that electronic spin fluctuations can lead to a reduction in the nuclear Schottky contribution [bertinEffectiveHyperfineTemperature2002]. The corresponding fluctuations in the hyperfine field can diminish the population differences between the 141Pr Zeeman levels, resulting in a suppressed , seen e.g. in Pr-based pyrochlores [maclaughlinUnstableSpiniceOrder2015, martinDisorderQuantumSpin2017a]. Another possibility would be a partial decoupling of the nuclear degrees of freedom due to a lack of phonons (or magnons) at these temperatures [abragamPrinciplesNuclearMagnetism, isonoSpinlatticeDecouplingTriangularlattice2018a]. The nuclear spin system only couples indirectly to the external heat bath through excitations of the lattice/electronic spins, so without an abundance of such low energy modes we cannot probe efficiently the nuclear heat capacity. A similar decoupling phenomenon was observed in the sister material Nd3BWO9 above the pseudospin saturation [flavianMagneticPhaseDiagram2023a].
In Pr3BWO9, the purported nuclear contribution is suppressed immediately below the onset temperature of the calculated , which heavily favours the latter scenario. For fast fluctuating electronic spins, the peak position should be preserved, while a decoupling of nuclear spins would become effective for , meaning that only the high temperature tail of the nuclear heat capacity can be observed. Towards lower fields, the hyperfine splitting decreases significantly, further suppressing until it is no longer visible below T.
@@ III.6 Ultrasound
Given the strong spin-orbit coupling inherent in rare earths, it is instructive to consider the effects of magnetoelastic coupling, as even small changes in the crystal structure can potentially have drastic consequences on the magnetism (and vice versa). To that end, we probe the ultrasound velocity in Pr3BWO9, focusing on the longitudinal mode propagating with km/s. Figure 7 displays the relative change in sound velocity as a function of magnetic field. In agreement with the heat capacity data, we do not detect any signs of phase transitions down to mK even for fields exceeding 40 T, suggesting that Pr3BWO9 retains a quantum disordered ground state in the entire phase space. Instead, we observe a small decrease in sound velocity at low fields (i.e. a small “softening” of the lattice), followed by a strong quasi-linear increase. Both the earlier crossover between these two regimes and the overall stronger magnetoelastic response for fields along are linked to the easy-axis nature of Pr3BWO9, explained mostly by a difference in -factors. Furthermore, qualitatively follows the single magnon gap (see Figure 6(b)), where the crossover appears when the crystal field splitting acts on the same scale as the Zeeman energy . In fact, these observations are not surprising, as the dominant contribution to the magnetoelastic energy density in rare earths typically originates from a linear single-ion strain interaction [luthiPhysicalAcousticsSolid1967]. Lattice vibrations modulate the crystal electric field, which directly determines the orbital state of each ion, resulting in a magnetoelastic response dictated by single-ion properties. Though the overall scale of of course depends on the field orientation, the sign of the effect does not. Because there is no transverse -factor, fields can only couple to the local pseudospin component, resulting in an analogous deformation to the charge density independent of field direction. We note that in principle, the phase sensitive detection method utilized here responds also to magnetostrictive changes in the sample length along the propagation direction, i.e. we measure the phase shift as denoted in Figure 7. However, barring any (magneto-)structural phase transitions or incipient soft modes thereof, the change in lattice spacing should be orders of magnitude weaker than the present effects and will therefore be ignored here.
@@
Perhaps more interesting is the significant temperature dependence shown in the inset of Figure 7. At zero field, there is a rapid increase in sound velocity upon heating, which becomes heavily suppressed at fields as low as T, gradually giving way to -independent behavior. All data are taken below , so effects of CEF population should be negligible and these features more likely result from a change in the spin correlations. But even though such a lattice “hardening” at zero field can be easily rationalized as a weakening of magnetic correlations, it is unclear how its sudden suppression at extremely low fields could be explained in this scenario. Indeed, there is no obvious energy scale associated with this suppression - both Weiss temperature and crystal field gap are significantly larger - and neither susceptibility nor magnetization exhibit significant features at such low fields, making this behavior somewhat anomalous.
All in all, magnetoelastic effects appear to be rather subdued in Pr3BWO9, inducing only gradual changes in sound velocity mostly accounted for by single-ion effects, seemingly proving unimportant to our understanding of the minimal magnetic model.
@@ III.7 Muon Spin Relaxation
As a complement to thermodynamic probes, the ground state magnetism in Pr3BWO9 is characterized by means of muon spin relaxation. The asymmetry spectra of the zero-field SR data at selected temperatures are displayed in Figure 8(a). Contrary to expectations, the low temperature data show a clear dip at s followed by the partial recovery of a weakly decaying tail, amounting to of the initial polarization. The signal remains temperature independent below K, whereas at higher temperatures the dip weakens and shifts to later times, gradually giving way to paramagnetic relaxation above K. A Kubo-Toyabe-like (KT) -tail provides strong evidence for quasi-static magnetism at the SR timescale [leeMuonScienceMuons2017], whereas the lack of further oscillations implies a frozen disordered state instead of collective long-range order.
@@
With this model, we can successfully account for the zero field relaxation in the entire temperature range 30 mK 200 K - the resulting fit parameters are provided in Figure 9. The width of the internal field distribution remains constant below 5 K before smoothly starting to decrease, only tending towards zero at the highest measured temperatures. Both the relaxation rate and the KT-exponent show a broad hump around 50 K, consistent with a slow crossover towards paramagnetic relaxation at high temperatures. Apart of that, the only prominent feature in the data is a step-like change seen around K, present in all fit parameters except . Especially the step in is directly visible in Figure 8(a) as a shift of the late-time asymmetry tail between the 1 K and 2.5 K curves. In fact, the same feature remains present in a T longitudinal field (see Figure 8(b)), even though the weak internal fields must be fully decoupled in this regime, suggesting a dynamical origin for the anomaly. We can explain this behavior by invoking the existence of several stopping sites 222In oxides, the muon tends to stop away from a nearby oxygen site, analogous to an OH bond [leeMuonScienceMuons2017]. Due to the complex low-symmetry crystal structure with 18 oxygen ions per unit cell, careful DFT calculations would be required to determine the potential candidate stopping sites. Such an analysis is beyond the scope of this work - it would not help in analyzing our SR results, as any model analogous to Eq. (6) with several stopping sites strongly over-parameterizes the data.: The residual dynamics in some fraction of the sample slows down at low temperatures until the tail relaxation of that component approaches zero within our time window. This brings about an increase in the “background” fraction , as well as a change in the effective relaxation rate . It is also consistent with the appearance of a stretching exponent for the dynamical relaxation, which is expected in the presence of several different relaxation times. The only magnetic energy scale roughly matching this K feature would be the that of the two-ion interactions meV, indicating that it could be related to the onset of some residual AFM correlations, which quench/slow down the dynamics in a fraction of the sample.
@@
In order to pinpoint the nature of the weak quasi-static magnetism, we compare the mT observed at low temperatures to typical internal field values expected in different scenarios. For instance, similar Pr-based materials that order magnetically typically exhibit a strong T dipolar field at the muon site [AnandPRB1072023], roughly two orders of magnitude larger than seen in experiment. This clearly contradicts the possibility of a glassy ground state, where the SR relaxation would be dominated by fully frozen electronic moments . Moreover, the internal field distribution is also atypical of nuclear magnetism (the 141Pr nuclei carry a spin ), which usually leads to a temperature-independent width in the range of mT - seemingly ruling out nuclear moments as a source of the KT-like behavior. But as we are dealing with a non-Kramers system endowed with strong hyperfine coupling and a low-lying CEF level, there is a well known mechanism for amplifying the effective nuclear moments, aptly named enhanced nuclear magnetism [bleaneyEnhancedNuclearMagnetism1973, blundellQuantumMuon2023]. Due to the large Van-Vleck susceptibility, the nuclear spins can partially polarize the electronic shell, giving rise to “dressed” nuclear moments
@@ III.8 Inelastic Neutron Scattering
| Distance (Å) | Value (meV) | Value (meV) | |
| - | 1.62(1) | 1.62(1) | |
| 3.96 | 0.30(1) | 0.0068(2) | |
| 3.96 | -0.11(1) | -0.0026(2) | |
| 4.26 | 0.06(1) | 0.0053(8) |
The inelastic neutron spectra collected at AMATERAS are summarized in Figure 10 and Figure 11. These are raw data without background subtraction, taken at mK. The spectrum in Pr3BWO9 seems to be composed of two kinds of excitations, featuring a coexistence of sharp spin-wave-like modes and broad continuum excitations.
@@
In order to analyze the observed spectrum, we fit a number of constant- cuts to a combination of one Gaussian-broadened (continuum) and several resolution-limited (spin waves) modes, extracting almost 600 dispersion points across the entire Brillouin zone. These data are analyzed by means of linear spin-wave theory (LSWT) calculations using the SUNNY software package [sunny_git]. We perform a global LSWT fit to the observed mode energies, based on the two-level TFIM Hamiltonian Eq. (4) derived above. In accordance with our thermodynamic data, the ground state is assumed to be quantum disordered in these calculations, corresponding to a fully polarized state in the pseudospin language of our transverse-field Ising model, where each spin points along its local -axis. Therefore, LSWT is expected to be an excellent approximation and should allow for an accurate estimation of the “true” (non-renormalized) exchange constants [coldeaDirectMeasurementSpin2002].
@@
The calculated intensities can be improved significantly by using the full CEF Hamiltonian
i.e. working with coherent states [dahlbomRenormalizedClassicalTheory2023] instead of dipoles polarized along the transverse field direction. Due to the mapping between the TFIM and CEF model, the exchange parameters in the latter framework can be obtained directly by rescaling those from the LSWT fit as
@@
where the matrix element accounts for the change in spin length. The geometrical factor reflects the fact that we are using Heisenberg couplings in the CEF model; projecting down to the low-energy degrees of freedom will anyways render void all exchange components not along the local Ising axes. We note that because of the much higher energy resolution on AMATERAS compared to the CEF spectra obtained on EIGER, the crystal field parameters in Table 4 are also rescaled slightly by a global factor to obtain satisfactory agreement. This allows us to fine tune the observed gap energy without changing the anisotropies refined in bulk. The appropriately rescaled parameters for the CEF model are listed in Table 3. They produce the exact same dispersion as the TFIM couplings, but give a quantitative agreement on the relative intensity distribution. The enhanced ratio in this framework simply reflects the difference in angle between the easy axes of first and second neighbors.
@@
Finally, we attempt to empirically model the broad continuum excitations. Since all spin-wave branches are accounted for across the Brillouin zone and the thermodynamic probes rule out the presence of further accidentally degenerate CEF modes, it is unclear how any scenario of fractionalized excitations could give rise to the continuum seen in Pr3BWO9. Instead, the fact that the latter seems to mimic the spin-wave dispersion favors a scenario of structural disorder. Due to the lack of time reversal symmetry in Pr3+, the splitting between our quasi-doublet states is highly sensitive to changes in the local crystal structure [martinDisorderQuantumSpin2017a, liSpinDynamicsGriffiths2021]. As such, any local distortion due to e.g. strains, impurities or domain walls will slightly alter the CEF scheme, resulting in a distribution of gaps . As we observe both broad and sharp contributions in our spectra, we have to assume that some ions are sufficiently removed from such impurity sites to give rise to the resolution limited spin-waves, while the remaining ones possess a random distribution of gaps, accounting for the continuum. An investigation of the local strain fields in Pr3BWO9 goes beyond the scope of this work, though we find that a simple Gaussian distribution is sufficient for modeling the spectrum.
@@
Indeed, we can successfully reproduce our inelastic data in this disorder scenario by adding a Gaussian-broadened copy of the calculated spin-wave spectrum with the appropriate spectral weight. To that end, we introduce three new global parameters: The fraction of affected magnetic sites, the Gaussian width meV reflecting the degree of randomness and the mean meV of the distribution (to obtain complete agreement with experiment we choose ). Presumably, quenched disorder would also induce some variation in the exchanges, but here we push all randomness onto the Gaussian width , in order not to over-parameterize our model. The calculated spectra obtained with these empirically optimized parameters in the CEF model framework are displayed in Figures 10(k,l,m,n,o) and 11(e,f,g,h). Our model can semi-quantitatively account for all experimental features, including the coexistence of resolution-limited and broadened modes, as well as the intensity distribution across the entire Brillouin zone.
@@
As a cross check, we calculate the single-ion heat capacity in the proposed disorder scenario, shown in Figure 5 as composed of Schottky terms with the distribution of gaps extracted from the neutron data. Unlike the “clean” single-ion calculation, this model closely approximates the measured curve in the entire temperature range with no free parameters, including the power-law like behavior at low temperatures. As seen in Figure 6, we can even reproduce the results at finite fields, although the small hump seen at low temperatures remains absent, consistent with our explanation in terms of partially decoupled nuclear degrees of freedom. This striking parameter-free agreement, together with the excellent fit of our inelastic neutron data, strongly justify our model assumptions, confirming that the magnetism of Pr3BWO9 should be understood in terms of dispersive crystal field excitons, with many sites affected by some form of structural disorder that locally tunes the average CEF scheme.